# Quasi-amorphous crystal

### v1

@@ Abstract
Brittle plastic yielding is a salient feature of well-annealed glassy materials. Here we show that the same behavior is characteristic of perfect crystals after they experience mechanically driven elastic instability leading to massive nucleation of dislocations. We argue that such ’preparation’ effectively converts an atomic configuration from crystalline to quasi-amorphous. To understand the nature of the subsequent mechanical response, which is reminiscent of quasi-brittle yielding we study an athermal model 2D crystal subjected to quasistatic loading. We show that the intermittent pre- and post-yield dislocation avalanches exhibit power law statistics with matching exponents. The computed value of these exponents is indicative of marginal stability.
The complexity of plastic behavior of crystalline solids is revealed by pre- and post-yield intermittent fluctuations accompanied by scale-free organization of dislocations  Zaiser (2006); Jian et al. (2024); Ortiz (1999); Papanikolaou et al. (2017); Kosterlitz (2016); Khantha and Vitek (1997); Hochrainer et al. (2014); Langer (2019); Derlet and Maaß (2016); Ispánovity et al. (2010). Such features are also typical of amorphous plasticity  Kamani and Rogers (2024); Berthier et al. (2024); Shang et al. (2020, 2024); Singh et al. (2020); Richard et al. (2021); Pollard and Fielding (2022); Lin et al. (2014); Parisi et al. (2017); Derlet and Maaß (2021). To build a conceptual link between the two phenomena, we study in this Letter a relatively transparent prototypical example of a 2D square crystal subjected to quasistatic loading in athermal conditions.
@@
To prepare a ’generic’ crystal we start with a pristine dislocation-free configuration and bring the associated perfect crystal to the point of mechanical instability which is resolved through massive dislocation nucleation. We argue that such ’preparation’ effectively converts an atomic configuration from crystalline to quasi-amorphous Lehtinen et al. (2016); Sethna et al. (2017); Ovaska et al. (2017). The subsequent plastic flow begins with a hardening stage, which involves a superposition of spatially localized dislocation rearrangements. This ’microplasticity’ regime is terminated abruptly with a quasi-brittle yield leading to the formation of multi-grain pattern involving global shear bands. The post-yield plastic flow is characterized by a stress plateau with superimposed fluctuations representing broadly distributed dislocation avalanches. Similar quasi-brittle yield has been observed in well-annealed glasses Nicolas et al. (2018); Rodney et al. (2011); Popović et al. (2018); Tanguy (2021); Talamali et al. (2012); Parley and Sollich (2024); Karmakar et al. (2010); Ozawa et al. (2018); Divoux et al. (2024). Despite the absence of dislocations in amorphous plasticity, behind both phenomena is an evolving elastic incompatibility Moshe et al. (2015); Baggioli et al. (2022); Bowick and Giomi (2009); Katanaev and Volovich (1992); Sozio and Yavari (2020) with underlying elementary mechanical events represented by shear eigenstrains Picard et al. (2004); Budrikis et al. (2017); Cai et al. (2006).
@@
To reveal the universal features of the criticality exhibited by crystal and amorphous plasticity, we perform a detailed study of dislocation dynamics in a prototypical athermal square crystal subjected to quasistatic loading by shear strain applied along one of the slip directions. Our main tool is a novel mesoscopic tensorial model (MTM) of crystal plasticity representing a conceptual trade-off between continuum and atomic descriptions Salman and Truskinovsky (2011, 2012); Baggio et al. (2019); Salman et al. (2021); Perchikov and Truskinovsky (2024); Baggio et al. (2023a, b); Zhang et al. (2020). Crucially, it allows one to capture the intermittent nature of dislocation avalanches while resolving topological transitions involved in nucleation and annihilation of dislocations without any specialized phenomenological assumptions Weiss et al. (2021). In contrast to other computational approaches to micro-scale plasticity  Arsenlis et al. (2007); Zepeda-Ruiz et al. (2021); Ponga et al. (2020); Lafourcade et al. (2025); Chan et al. (2010); Salvalaglio et al. (2020); Zhong et al. (2025); Javanbakht and Levitas (2016); Beyerlein and Hunter (2016); Koslowski et al. (2002); Wang et al. (2003); Devincre et al. (2008); Ispánovity et al. (2014); Gómez-García et al. (2006); Po et al. (2014); Lu et al. (2022); Bertin et al. (2024); Nasir Tak and Guruprasad (2025); Starkey et al. (2020); Zhang et al. (2025); Acharya (2022); Acharya et al. (2006); Rodney and Phillips (1999); Van Koten and Luskin (2011); Miller and Tadmor (2002); Zhang et al. (2018) the MTM allows one to deal with statistically meaningful number of dislocations while accounting adequately for large lattice rotations and presenting geometrically exact description of lattice-invariant shears Engel (1986); Michel (2001); Schwarzenberger (1972); Pitteri and Zanzotto (2002).
@@
The main idea of the MTM approach is to associate with material points an effective energy landscape whose tensorial periodicity accounts for crystallographically-invariant deformations. The resulting Landau-type model is characterized by infinitely many equivalent energy wells Ericksen (1970, 1973, 1977, 1980, 1989); Parry (1976, 1977, 1998); Folkins (1991); Conti and Zanzotto (2004); Bhattacharya et al. (2004); Salman and Truskinovsky (2011, 2012); Baggio et al. (2019); Salman et al. (2021); Perchikov and Truskinovsky (2024); Baggio et al. (2023a, b); Zhang et al. (2020) and plastically deformed crystals emerge as a multi-phase mixtures with dislocations playing the role of effective domain boundaries. The small scale regularization is achieved by mesoscopic discretization. Despite operating with engineering concepts of stress and strain the resulting computational code correctly captures not only long-range elastic interactions of dislocations but also describes adequately their short-range interactions involved in the formation of complex dislocational entanglements. Some elements of the MTM approach can be found in Denoual et al. (2010); Gao et al. (2019, 2020a); Gao (2020); Gao et al. (2020b, c); Biscari et al. (2015), see also some parallel recent developments in Arbib et al. (2020, 2023, 2025).
Consider the continuum deformation field $\bf y=\bf y(\bf x),$ where $\bf y$ and $\bf x$ are positions of material points in the current and reference configurations, respectively. Due to frame indifference constraint, the strain energy density $\varphi$ can depend on the deformation gradient $\bf F=\nabla\bf y,$ only through the metric tensor ${\bf C}={\bf F}^{T}{\bf F}$ which then plays the role of a tensorial order parameter. The function $\varphi(\bf C)$ should respect the symmetry of an underlying Bravais lattice extended beyond the conventional point group. For instance, to capture lattice invariant shears in 2D theory we should require that
$\varphi({\bf C})=\varphi({\bf m}^{T}{\bf C}{\bf m}),$ (1)
with the unimodular integer valued matrix $\bf m$ belonging to $GL(2,\mathbb{Z})$, see SM for details. In view of this symmetry, the surface $\det{\bf C}=1$ in the 3D space $(C_{11},C_{22},C_{12})$ is tessellated into an infinite number of equivalent periodicity domains; if the energy density is known in one of such domains, it can be extended to the whole configurational space by $GL(2,\mathbb{Z})$ symmetry. To visualize this tessellation, it is convenient to project stereographically the infinite surface $\det{\bf C}=1$ on a unit (Poincaré) disk, see Fig. 1 where we show in gray the fundamental (minimal periodicity) domain $\mathscr{D}$.
In our numerical experiments we used the polynomial function $\varphi(\bf C)$ proposed in Conti and Zanzotto (2004) which is the simplest to ensure stress continuity, see SM for details. In Fig. 2 we illustrate the corresponding multi-well energy landscape with equivalent energy wells representing replicas of the reference square lattice. One can see that such a landscape contains a network of low energy valleys depicting in our case simple shears parallel to crystallographic slip planes and therefore mimicking the conventional plastic ’mechanisms’ operating in square crystals.
@@
Figure 1: $GL(2,\mathbb{Z})$-induced tessellation of the configuration space of metric tensors $C_{11},C_{22},C_{12}$ with $\det\mathbf{C}=1$ stereographically projected on the Poincaré disk $p^{2}+q^{2}<1$ where $p=t(C_{11}-C_{22})/2$, $q=tC_{12}$, and $t=2/(2+C_{11}+C_{22})$. The highlighted domain $\mathscr{D}$ is the fundamental (minimal periodicity) domain.
@@
The presence of the energy barriers separating individual energy wells makes the function $\varphi(\mathbf{C})$ highly nonconvex, see Fig. 2. Therefore, the corresponding scale-free Landau model is degenerate and needs to be regularized through the introduction of a cut-off length scale. This is achieved in the MTM by projecting the continuum problem on a uniform mesoscale grid and associating the periodic elastic energy with discrete elements whose linear size $h$ becomes a physical parameter. In this way the infinite dimensional space of continuum deformation fields is reduced to a finite dimensional set of compatible, piece-wise affine mappings, see SM for details.
Figure 2: Energy landscape in the space of metric tensors (Poincaré) disk shown in Fig. 1. Color indicates the energy level with only the lowest energy levels shown. The deepest energy wells correspond to equivalent representations of the same square lattice.
The driven system of this type will experience a rich repertoire of instabilities. Specificaly, an incremental elastic energy minimization of a system subjected to quasistatic loading will generically lead to repeated jump discontinuities representing branch switching events. The energy losses associated with such jumps will represent plastic dissipation; in the continuum limit the jumps will merge giving rise to rate-independent plasticity Puglisi and Truskinovsky (2005).
Figure 3: Dislocated configuration emerging after a homogeneous lattice is mechanically driven to the threshold of elastic instability. Colors indicate the level of the shear component of the Cauchy stress tensor $\sigma_{xy}$. The highest stress levels of both signs correspond to the location of dislocation cores.
In our numerical experiments a square crystal was represented by $N\times N$ nodes, with $N=400$. The nodes were arranged into triangular finite elements. We loaded the corresponding athermal system quasi-statically by applying the affine displacement field
${\bf u}(\alpha,{\bf x})=(\bar{\bf F}(\alpha)-\mathbb{1}){\bf x},$ (2)
with
$\bar{\bf F}(\alpha)={\bf I}+\alpha{\bf e}_{1}\otimes{\bf e}_{2},$
@@
where ${\bf e}_{i}$, $i=1,2$, is the orthonormal basis of the reference square lattice. In this way we imposed simple shear along one of the principal slip directions with shear amplitude $\alpha$ serving as the loading parameter. By changing this parameter in strain increments of order $10^{-6}$, we advanced the loading. After each such increment the displacement field was updated by the energy minimization algorithm. Details of the discretization, of the incremental energy minimization algorithm and of the implementation of the loading protocol under periodic boundary conditions can be found in SM .
Figure 4: (a) Stress-strain response of a quasi-amorphous crystal shown in Fig. 3. (b) Dislocated configuration in a typical pre-yield state characterizing the stage of ’microplasticity’. (c,d) Dislocated configuration characterizing different stages of unfolding of the quasi-brittle system size event. Colors indicate the level of the shear component of the Cauchy stress tensor $\sigma_{xy}$.
@@
To obtain a ’generic’ sample, we first loaded a pristine (defect-free) crystal till the affine configuration became elastically unstable. The corresponding critical value of the loading parameter $\alpha=\alpha_{c}\approx 0.138$ is very close to the threshold of positive definiteness of the acoustic tensor Simpson and Spector (1989); Grabovsky and Truskinovsky (2013); Steigmann et al. (2023), see the explicit formulas in SM . The breakdown of an elastic state takes the form of a major system-size inelastic event. It leads to multiple nucleation of dislocations which apparently transforms a perfect crystal into a quasi-amorphous system, see Fig. 3. We observe that the unfolding of the elastic instability is arrested at a state which one can expect to be only marginally stable.
@@
We continued to load the obtained ’generic’ sample using the same athermal quasistatic protocol and the ensuing stress-strain response is shown in Fig. 4(a). Its salient feature is the presence of irregularly placed elastic branches interrupted by intermittent stress drops; the latter represent plastic avalanches involving partial unlocking of dislocation structures. Among these avalanche-type events, one major stress drop stands out and can be interpreted as a plastic yield separating two different regimes. While the first (pre-yield) regime is characterized by quasi-elastic hardening, the second (post-yield) regime is represented by a stress plateau which suggests quasi-stationary plastic flow, see Fig. 4(a).
@@
Our Fig. 4(b) illustrates the stress field in a typical pre-yield state $A$. The restructuring of the dislocation pattern is relatively minor with dislocations moving only between the preexisting locking sites, cf. Fig. 3. Such response is known to be a characteristic pre-yield feature for both amorphous and (defective) crystalline solids Maaß and Derlet (2018); Papanikolaou et al. (2017); Sparks and Maaß (2018); Sparks et al. (2019); Rizzardi et al. (2022); Duan et al. (2023). This ’microplasticity’ regime ends with the system-size event reminiscent of quasi-brittle plastic yield encountered in well-annealed glasses and also in some sub-micron crystals Bei et al. (2008); Chrobak et al. (2011); Wang et al. (2012); Cui et al. (2017); Greer and Hosson (2011); Brenner (1956, 1958); Sharma et al. (2018); Mordehai et al. (2018); Lilleodden and Nix (2006); Corcoran et al. (1997). As we see from Fig. 4(a) it takes the form of a two-stage transition $A\to B\to C$; the snapshots of the spatial stress configurations in the states $B$ and $C$ are shown in Fig. 4(c,d). The associated global restructuring reveals collective dislocation activity manifested through the emergence of system-spanning shear bands.
@@
Additional details can be seen in Fig. 5 where we illustrate the strain-energy density field in the state $C$. We see the development of low-energy rotated patches of the original lattice forming intricate polycrystalline grain texture with misorientation of neighboring grains controlled by the overall compatibility of the deformation field as explained in Baggio et al. (2023b); the elastic energy is localized around dislocation-rich grain boundaries. In the inset in Fig. 5 we made dislocations of different sign clearly visible: they were identified via Delaunay triangulation and nodes with five/seven neighbors are shown in blue/red. One can see the emergence of energy minimizing structure of grain boundaries Sutton (1984); Priester (2012); Randle (2024) with ubiquitous presence of $\Sigma 5$ grain boundaries, see Baggio et al. (2023b) for the explanation.
@@
Figure 5: The micro-configuration configuration corresponding to the state $C$ marked in Fig. 4. Colors indicate the magnitude of strain-energy density. The inset shows a single grain and details the dislocation structure of the grain boundaries.
@@
The post-yield regime, characterized by relative stabilization of the average stress level, can be interpreted as a quasi-stationary non-equilibrium steady state (NESS). A representative snapshot of the micro-configuration of the crystal in the state $D$ is shown in Fig. 6 where we see further development of the polycrystalline grain texture. During intermittent avalanches characterizing this regime larger grains occasionally merge while small grains continue to emerge. In this way the self-organization of dislocations continues to generate additional scales which contribute to the formation of the global hierarchical structure. Through the transitions from isolated dislocation motion to collective dislocation behavior the system acquires access to a much broader repertoire of relaxation mechanisms. A detailed analysis of the avalanche structure shows that while in the pre-yield regime plastification takes place primarily in the form of isolated, linear arrangements of transformed elements, a typical post-yield avalanche reveals extended plastified regions with complex branching.
Figure 6: Mature shear bands in the regime of stationary post-yield plastic flow represented by the state $D$ marked in Fig. 4. Colors indicate the level of the shear component of the Cauchy stress tensor $\sigma_{xy}$.
An important window into the micro-mechanics of solids undergoing plastic flow is provided by the study of the statistical structure of plastic fluctuations Papanikolaou et al. (2017); Weiss et al. (2015); Zaiser (2006); Tarp et al. (2014); Miguel and Zapperi (2006); Alava et al. (2014); Sethna et al. (2017). Thus, it is known that in amorphous solids the probability distribution of the energy drops associated with individual avalanches $\Delta W=W_{\text{before}}-W_{\text{after}}$, where $W$ is the total elastic energy of the crystal defined in SM , is represented by power laws of the form
$P(x)\sim x^{-\epsilon}e^{-\lambda x}$ (3)
@@
with the same exponent $\epsilon$ but different values of the cut-off parameter $\lambda$ for pre-yield and post-yield regimes Tyukodi et al. (2016); Budrikis et al. (2017); Lerner et al. (2018); Ozawa et al. (2022). Our results, illustrated in Fig. 7, show that for ’generic’ crystals prepared using our protocol, the statistics of avalanches follows the same pattern. In particular, similar to the case of well-annealed glasses Papanikolaou et al. (2017); Weiss et al. (2021); Sethna et al. (2001); Ispánovity et al. (2014); Alava et al. (2014); Lehtinen et al. (2016), the scaling behavior remains basically unchanged across the quasi-brittle yielding transition. The obtained power law range extends over up to six decades with pre-yield and post-yield energy distributions sharing the same exponent $\epsilon=1.0$. The invariance of the exponent $\epsilon$ across the yielding transition suggests that while the spatial organization of plastic events changes drastically from local and non-cooperative to global and collective, the statistical distribution of events remains insensitive to these details responding only to the global organization of the underlying energy landscape Shang et al. (2020); Oyama et al. (2021).
Figure 7: Probability distribution of energy drops during avalanches in pre- and post-yield regimes.

### v2

# Self-induced marginality in plastically deformed crystals

@@ Abstract
Quasi-brittle plastic yielding is a salient feature of well-annealed glassy materials. Here we show that the same behavior is characteristic of perfect crystals after they experience mechanically driven elastic instability leading to massive nucleation of dislocations. We argue that such ’preparation’ effectively converts an atomic configuration from crystalline to quasi-amorphous. To understand the nature of the subsequent intermittent mechanical response we study a model 2D crystal subjected to AQS driving and show that both pre- and post-yield dislocation avalanches exhibit power law statistics with similar exponents indicative of self-induced marginal stability.
The complexity of plastic response of strain-driven crystalline solids is revealed by intermittent stress fluctuations accompanied by scale-free spatial organization of dislocations  Papanikolaou et al. (2017); Hochrainer et al. (2014); Derlet and Maaß (2016); Ispánovity et al. (2010). The same general features are also typical of amorphous plasticity  Berthier et al. (2024); Shang et al. (2020); Singh et al. (2020); Lin et al. (2014). To build a conceptual link between the two phenomena, we study in this Letter a relatively transparent prototypical example of a 2D square crystal subjected to athermal quasistatic (AQS) proportional loading by shear strain applied along one of the slip directions. To prepare a ’generic’ crystal we start from a pristine dislocation-free crystal and shear it it to the point of mechanical instability which is resolved through massive dislocation nucleation. We argue that such ’preparation’ effectively converts an atomic configuration from crystalline to quasi-amorphous, e.g. Ovaska et al. (2017). This term is used to refer to a solid that is structurally crystalline but mechanically amorphous in the sense that its mechanical response is dominated by non-affine relaxations, e.g. Zaccone and Scossa-Romano (2011).
@@
We observe that plastic flow of such quasi-amorphous crystals begins with a hardening stage involving small spatially localized dislocation rearrangements. This ’microplasticity’ regime is terminated abruptly with a quasi-brittle yield leading to the formation of multi-grain pattern dominated by global shear bands. The post-yield plastic flow is characterized by a stress plateau with superimposed fluctuations representing broadly distributed dislocation avalanches. Similar quasi-brittle yield has been observed in well-annealed glasses Tanguy (2021); Parley and Sollich (2024); Ozawa et al. (2018). Despite the absence of dislocations in amorphous plasticity, both phenomena can be viewed as representing the evolution of elastic incompatibility with underlying elementary mechanical events represented by shear eigenstrains Picard et al. (2004); Budrikis et al. (2017); Baggioli et al. (2022).
@@
Our main tool is a novel mesoscopic tensorial model (MTM) of crystal plasticity representing a conceptual trade-off between continuum and atomic descriptions Salman and Truskinovsky (2011, 2012); Baggio et al. (2019); Salman et al. (2021); Perchikov and Truskinovsky (2024); Baggio et al. (2023a, b); Zhang et al. (2020). Crucially, it allows one to capture the intermittent nature of dislocation avalanches while resolving topological (connectivity) restructuring involved in nucleation and annihilation of dislocations without any specialized phenomenological assumptions Weiss et al. (2021). In contrast to other computational approaches to micro-scale plasticity, e.g. Zepeda-Ruiz et al. (2021); Chan et al. (2010); Salvalaglio et al. (2020); Javanbakht and Levitas (2016); Beyerlein and Hunter (2016); Koslowski et al. (2002); Wang et al. (2003); Devincre et al. (2008); Ispánovity et al. (2014); Gómez-García et al. (2006); Po et al. (2014); Starkey et al. (2020); Zhang et al. (2025), the MTM allows one to deal with statistically meaningful number of dislocations while accounting in a geometrically exact way for finite deformations  Engel (1986); Pitteri and Zanzotto (2002).
@@
The main idea of the MTM approach is to associate with material points an effective energy landscape whose tensorial periodicity accounts for crystallographically-invariant deformations. The resulting Landau-type model is characterized by infinitely many equivalent energy wells Ericksen (1970); Parry (1998); Folkins (1991); Conti and Zanzotto (2004); Bhattacharya et al. (2004); Salman and Truskinovsky (2011); Baggio et al. (2019); Perchikov and Truskinovsky (2024) and plastically deformed crystals emerge as a multi-phase mixtures with dislocations playing the role of effective domain boundaries. The small scale regularization is achieved by mesoscopic discretization. Despite operating with engineering concepts of stress and strain the resulting computational code correctly captures not only long-range elastic interactions of dislocations but also describes adequately their short-range interactions involved in the formation of complex dislocational entanglements, see also variations on the same theme in Denoual et al. (2010); Gao et al. (2019); Biscari et al. (2015); Arbib et al. (2023). We also note that the 2D version of the MTM, used in our numerical experiments, has some formal similarities with a single-slip-plane phase field model of Peierls-Nabarro type, which was found fully adequate to represent the critical nature of plastic flows, e.g. Koslowski et al. (2002, 2004). The main difference of our approach is the use of finite strain bulk elasticity compatible with multi-slip and capturing adequately lattice rotations which produced different values of critical exponents revealing the marginal stability of the underlying configurations.
Refer to caption Refer to caption
(a) (b)
@@
Figure 1: (a) $GL(2,\mathbb{Z})$-induced tessellation of the configuration space of metric tensors $C_{11},C_{22},C_{12}$ with $\det\mathbf{C}=1$ stereographically projected on the Poincaré disk $p^{2}+q^{2}<1$ where $p=t(C_{11}-C_{22})/2$, $q=tC_{12}$, and $t=2/(2+C_{11}+C_{22})$. The highlighted domain $\mathscr{D}$ is the fundamental (minimal periodicity) domain. (b) Representative energy landscape with colors indicating the energy level. Only the lowest energy levels are visible with the bottoms of the wells corresponding to equivalent replicas of the reference square lattice.
Consider the continuum deformation field $\bf y=\bf y(\bf x),$ where $\bf y$ and $\bf x$ are positions of material points in the current and reference configurations, respectively. Due to frame indifference constraint, the strain energy density $\varphi$ can depend on the deformation gradient $\bf F=\nabla\bf y$ only through the metric tensor ${\bf C}={\bf F}^{T}{\bf F}$ which then plays the role of a tensorial order parameter. In MTM we assume that the function $\varphi(\bf C)$ additionally respects the symmetry of an underlying Bravais lattice. Such symmetry is assumed to be global extending beyond the conventional point group and capturing lattice invariant shears. In 2D it reduces to the constraint that $\varphi({\bf C})=\varphi({\bf m}^{T}{\bf C}{\bf m}),$ where unimodular integer valued matrices $\bf m$ belong to $GL(2,\mathbb{Z})$ SM . In view of this symmetry, the surface $\det{\bf C}=1$ in the 3D space $(C_{11},C_{22},C_{12})$ is tessellated into an infinite number of equivalent periodicity domains; if the energy density $\varphi({\bf C})$ is known in one of such domains, it can be extended to the whole configurational space of tensors $\bf C$ by the symmetry. To visualize the implied tessellation, it is convenient to project stereographically the infinite surface $\det{\bf C}=1$ on a unit (Poincaré) disk, see Fig. 1(a). In our numerical experiments we used the simplest piece-wise polynomial function $\varphi(\bf C)$ proposed in Conti and Zanzotto (2004), see SM for details. In Fig. 1(b) we illustrate the corresponding multi-well energy landscape. Here the equivalent energy wells, which are somewhat distorted due to strereographic projection, represent the equivalent replicas of the reference square lattice. A characteristic feature of such a landscape is a network of low-energy valleys corresponding in our case to simple shears parallel to crystallographic slip planes. These valleys represent conventional plastic ’mechanisms’ and are assumed to be flat in the classical continuum crystal plasticity.
@@
Since the MTM preserves the energy barriers separating individual wells along the low-energy valleys, the function $\varphi(\mathbf{C})$ ends up being highly nonconvex. The corresponding continuum scale-free Landau model is then degenerate and needs to be regularized through the introduction of a cut-off length scale. This is achieved through the projection of the continuum problem on a uniform mesoscale grid while associating the nonconvex elastic energy to finite discrete elements. In other words, the infinite dimensional space of continuum deformation fields is reduced to a finite dimensional set of compatible, piece-wise affine mappings, with the linear scale of an individual finite element becoming a physical parameter SM . An incremental elastic energy minimization in a system of this type subjected to quasistatic loading will generically lead to a rich repertoire of instabilities with jump discontinuities representing elastic branch switching events. The energy losses associated with such jumps contribute to plastic dissipation; in the continuum limit the jumps will merge giving rise to rate-independent plasticity Puglisi and Truskinovsky (2005).
@@
In our numerical experiments a square crystal was represented by $N\times N$ nodes, with $N=100,200,400$. The nodes were arranged into triangular finite elements. We loaded the system quasi-statically by applying the affine displacement field ${\bf u}(\alpha,{\bf x})=(\bar{\bf F}(\alpha)-\mathbb{1}){\bf x},$ with $\bar{\bf F}(\alpha)={\bf I}+\alpha{\bf e}_{1}\otimes{\bf e}_{2},$ where ${\bf e}_{i}$, $i=1,2$, is the orthonormal basis of the reference square lattice. In this way we imposed simple shear along one of the principal slip directions with shear amplitude $\alpha$ serving as the loading parameter. The loading was advance in strain increments of order $10^{-6}$ and after each increment the displacement field was updated by the energy minimization algorithm. Details of the discretization, of the incremental energy minimization algorithm and of the implementation of the loading protocol under periodic boundary conditions can be found in SM .
Figure 2: Dislocated configuration emerging after a homogeneous lattice is mechanically driven to the threshold of elastic instability. Colors indicate the level of the $\sigma_{xy}$ component of the Cauchy stress tensor. The highest stress levels of both signs correspond to the location of dislocation cores. Here $N=400$.
@@
To obtain a ’generic’ sample, we first loaded a pristine (defect-free) crystal till the point of elastic instability of the affine configuration. The recorded critical value of the loading parameter $\alpha=\alpha_{c}\approx 0.138$ was found to be very close to the threshold of positive definiteness of the acoustic tensor Grabovsky and Truskinovsky (2013), see the explicit formulas in SM . To trigger the a generic instability we imposed a small Gaussian quenched disorder which was subsequently removed. The breakdown of an elastic state took the form of a massive nucleation of dislocations resulting in a highly inhomogeneous configuration of the type shown in Fig. 2. It is natural to assume that the unfolding of the elastic instability was arrested at a state which is only marginally stable and is therefore far from being random.
Figure 3: (a) Stress-strain responses of four different ’generic’ quasi-amorphous crystals of the type shown in Fig. 2; solid black curve indicates the average response. The stress-strain curves do not start from the origin because the global instability does not relaxes stresses to zero due to the emergence of self-equilibrated (residual) dislocation configuration. (b) Dislocated configuration in a typical pre-yield state characterizing the stage of ’microplasticity’. (c,d) Dislocated configuration characterizing different stages of unfolding of the quasi-brittle system size event. Colors indicate the level of the $\sigma_{xy}$ component of the Cauchy stress tensor. In (a-d) N=400.
@@
To study the mechanical response of the obtained ’generic’ sample we continued to load it quasistatically while recording the ensuing stress-strain relation. The results are shown in Fig. 3(a) for four different realization of disorder with the solid black curve indicating the average response. Its salient feature is the presence of irregularly placed elastic branches interrupted by intermittent stress drops; the latter represent plastic avalanches involving partial unlocking of dislocation structures with superimposed distributed nucleation/anihilation of dislocations. A major diffuse stress drop, which we interpret as the indication of the beginning of plastic yield, separates two markedly different regimes. While the quasi-elastic pre-yield regime is characterized by finite overall rigidity, the post-yield regime is represented by an effective stress plateau; in both regimes we observe intense plastic fluctuations, see Fig. 3(a).
@@
Our Fig. 3(b) illustrates the stress field in a typical pre-yield state $A$. Comparing to Fig.  2, the overall stress level is higher but the restructuring of the dislocation pattern is relatively minor with dislocations mostly displaced between the preexisting locking sites. Such ’microplasticity’ response is a characteristic pre-yield feature for both amorphous and (defective) crystalline solids Maaß and Derlet (2018); Papanikolaou et al. (2017); Sparks et al. (2019); Rizzardi et al. (2022) This regime ends with the system-size event which takes the form of single quasi-brittle plastic stress drop in well-annealed glasses and some sub-micron crystals Cui et al. (2017); Greer and Hosson (2011); Sharma et al. (2018); Mordehai et al. (2018). In Fig. 3(c,d) we show the snapshots of the spatial stress configurations in the intermediate state $B$ and the final state $C$ ending the transition to post-yield regime. The comparison of Fig. 3(b-d) shows that the associated global restructuring involves collective dislocation activity leading to the formation of system-spanning shear bands.
@@
Additional aspects of the state $C$ can be seen in Fig. 4 where instead of the stress field, as in our Fig. 3(d), we now show the strain-energy density field. In this representation the locations of individual dislocations become visible. The general pattern is the development of low-energy patches of the original lattice forming polycrystalline grain texture with elastic energy localized around dislocation-rich grain boundaries. In the inset presented in Fig. 4 we show the zoom in on the configuration of elastic elements; individual dislocations were identified via Delaunay triangulation and the nodes with five/seven neighbors were shown in blue/red. In particular, we see the ubiquitous presence of $\Sigma 5$ grain boundaries which can be explained by the fact that the misorientation of neighboring grains is controlled by the overall compatibility of the deformation field Baggio et al. (2023b).
@@
Figure 4: The micro-configuration configuration corresponding to the state $C$ marked in Fig. 3. Colors indicate the magnitude of strain-energy density. The inset shows a single grain and details the dislocation structure of the grain boundaries. Here N=400.
@@
The ensuing post-yield regime, characterized by the relative stabilization of the average stress level, can be broadly interpreted as a quasi-stationary non-equilibrium steady state. A representative snapshot of the micro-configuration of the crystal in the state $D$ is shown in Fig. 5(a) where we see progressive maturation of the polycrystalline grain texture. During intermittent avalanches characterizing this regime larger grains occasionally merge while smaller grains continue to emerge. In this way the self-organization of dislocations continues to generate additional scales which contribute to the formation of the global hierarchical structure. The complexity of the emerging configuration is illustrated in Fig. 5(b) where we show the corresponding strain distribution in the Poincaré disk. While in the reference state (pristine crystal) all configurational points were concentrated in the origin, now they are spread all over the configurational space with expected accumulation near the bottoms of at least five energy wells representing different replicas of the unstressed square lattice; for comparison, in the pre-yield regime the system explores at most three energy wells immediately adjacent to the reference state. In other words, plastic yield not only marks the transitions from isolated dislocation motion to collective dislocation behavior, but also implies the access to a much broader repertoire of relaxation mechanisms. For instance, a more detailed analysis of the avalanche structure illustrated in SM shows that while in the pre-yield regime plastification takes place primarily in the form of isolated, linear arrangements of transformed elements, a typical post-yield avalanche reveals extended plastified regions with complex branching.
Refer to caption Refer to caption
(a) (b)
Figure 5: (a) Shear band development in the regime of stationary post-yield plastic flow represented by the state $D$ in Fig. 3. Here N=400. Colors indicate the level of the shear component of the Cauchy stress tensor $\sigma_{xy}$. (b) Corresponding distribution of local values of the metric tensors in the configurational space projected on the Poincaré disk. Background color represents the energy landscape whose 3D representation was shown in Fig. 1(b). The color of the gray dots represents the number of the finite elements sharing the same metric tensor.
An important window into the micro-mechanics of solids undergoing plastic flow is provided by the study of the statistical structure of plastic fluctuations Papanikolaou et al. (2017); Weiss et al. (2015); Alava et al. (2014). In our case the relevant observables would be the energy dissipated at a single avalanche $\Delta W$ and the corresponding macroscopic stress drop $\Delta\Sigma=A\Delta\sigma_{xy}$, where $A$ is the surface area of the sample. It is known that in amorphous solids the probability distribution of these observables is compatible with the finite-size scaling ansatz $P(s;L)\;=\;s^{-x}\,\mathcal{F}\!\left(\frac{s}{L^{D_{x}}}\right),$ where $L=Nh$ is the linear system size, $h$ is the size of an element, $x$ is the power-law exponent with $x=\tau$ for stress drops and $x=\epsilon$ for dissipated energy and $D_{x}$ is the corresponding fractal dimension ($D_{\epsilon}$ or $D_{\tau}$, respectively). Finally $\mathcal{F}(u)$ is a universal scaling function. It is also known that in amorphous plasticity the exponents $\epsilon$ and $\tau$ are similar for pre-yield and post-yield regimes while the exponents $D_{\epsilon}$ and $D_{\tau}$ may be different Tyukodi et al. (2016); Budrikis et al. (2017); Lerner et al. (2018); Ozawa et al. (2022).
To determine these power-law exponents for our numerical experiments we employed a robust data-collapse optimization procedure detailed in SM . The resulting data collapse is illustrated in Fig.  6. One can see that for our ’generic’ quasi-amorphous crystals the statistics of avalanches follows the same pattern as for their amorphous counterparts. In particular, similar to the case of well-annealed glasses Papanikolaou et al. (2017); Weiss et al. (2021); Sethna et al. (2001); Ispánovity et al. (2014); Alava et al. (2014); Lehtinen et al. (2016), the scaling behavior remains basically unchanged across the quasi-brittle yielding transition.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 6: Optimal data collapses of the probability distributions onto universal master curves. (a, b) Collapse of the energy drops $\Delta W$ in the pre-yield generating exponents $\epsilon=0.9195$ and $D_{\epsilon}=0.3640$ (a) and post-yield regimes generating exponents $\epsilon=1.0207$ and $D_{\epsilon}=0.9836$ (b). Collapse of the macroscopic stress drops $\Delta\Sigma$ in the pre-yield generating exponents $\tau=0.9070$ and $D_{\tau}=0.3987$ (c) and post-yield regimes, generating exponents $\tau=1.2476$ and $D_{\tau}=1.4379$ (d).
@@
More specifically, the obtained power law range for energy drops extends over up to six decades with pre-yield and post-yield energy distributions sharing almost the same value of the exponent $\epsilon=1.0$. The invariance of this exponent across the yielding transition suggests that while the spatial organization of plastic events changes drastically from local and non-cooperative to global and collective, the statistical distribution of energy drops remains insensitive to these details responding only to the global organization of the underlying energy landscape Shang et al. (2020); Oyama et al. (2021).